OpenStudy (anonymous):

Which equation would best help solve the following problem? Adina has a rectangular garden that measures 9 m wide by 13 m long. She wants to increase the area to 192 m2 by increasing the width and length by the same amount. What will be the dimensions of the new garden? a. 192 = (9)(13) b. 192 = (9)(13 + x2) c. 192 = (9 – x)(13 – x) d. 192 = (9 + x)(13 + x)

OpenStudy (brunsonni):

D I think

OpenStudy (anonymous):

lets break this down into parts: the equation for Area of a rectangle is: A=L*W We know the Length is L = 13 and the Width is W = 9 So as is, the area is A=13*9=117 m^2 We want to add a bit to each side. Lets call that unknown amount "x" We end up with something like whats shown below:  Can you write an equation for the new length and the new width?